### v4

# On the Distributed Evaluation of Generative Models

Abstract
The evaluation of deep generative models has been extensively studied in the centralized setting, where the reference data are drawn from a single probability distribution. On the other hand, several applications of generative models concern distributed settings, e.g. the federated learning setting, where the reference data for conducting evaluation are provided by several clients in a network. In this paper, we study the evaluation of generative models in such distributed contexts with potentially heterogeneous data distributions across clients. We focus on the widely-used distance-based evaluation metrics, Fréchet Inception Distance (FID) and Kernel Inception Distance (KID). In the case of KID metric, we prove that scoring a group of generative models using the clients’ averaged KID score will result in the same ranking as that of a centralized KID evaluation over a collective reference set containing all the clients’ data. In contrast, we show the same result does not apply to the FID-based evaluation. We provide examples in which two generative models are assigned the same FID score by each client in a distributed setting, while the centralized FID scores of the two models are significantly different. We perform several numerical experiments on standard image datasets and generative models to support our theoretical results on the distributed evaluation of generative models using FID and KID scores.
1 Introduction
Deep generative models including variational autoencoders [1], generative adversarial networks (GANs) [2], and diffusion models [3] have attained remarkable results over a wide array of machine learning tasks. These frameworks’ success can be attributed to the large capacity of multi-layer neural networks in modeling complex distributions of image and text data as well as the intricate design of the training mechanisms in these algorithms. The promising results of these generative modeling schemes have inspired the development of several assessment methods to properly compare and rank a group of generative models using their produced data.
A standard approach to evaluating generative models is to compare the distribution of their generated data with the distribution of samples in a reference set, typically chosen to be the test sample set in a learning problem. This comparison is often made using a distance measure between the generative and reference distributions. Two standard distance-based metrics are Fréchet Inception Distance (FID) [4] and Kernel Inception Distance (KID) [5]. In a standard centralized comparison of generative models, only a single reference distribution is considered for the evaluation. Therefore, to rank a group of generative models in a centralized setting, one can sort their distances to the reference distribution.
On the other hand, several modern applications of deep generative models concern distributed settings where the reference data are provided by multiple clients in a network. A well-known distributed scenario is the federated learning setting [6], where several clients are connected to a server and aim to train and assess a decentralized model through their communications with the server node while preserving the privacy of their collected data. An important consideration for conducting evaluation in distributed settings is the potential heterogeneous data distributions across clients, since the background features of every client could lead to a different data distribution. For example, as displayed in Figure 1, the distribution of photos of clients in a mobile network could be highly diverse as they could possess different interests and characteristics.
Figure 1: Centralized (score-all) vs. distributed (score-avg) evaluation of generative models. The centralized and distributed FID-based evaluations can lead to different rankings, while we prove the KID-based evaluations result in the same ranking of a group of generative models.
In this work, we focus on evaluating generative models in distributed settings and attempt to study the assessment task’s complexity under diverse client data distributions. Following the existing literature on training generative models in distributed settings [7, 8, 9, 10, 11, 12], the ultimate goal is to perform a centralized evaluation where the reference set is the collective dataset of all clients’ data. In our analysis, we call the evaluation score following such a centralized approach the score-all metric, which means the score is computed considering a reference set with all the clients’ data. However, the commonly considered privacy constraints in distributed settings restrict the use of such an evaluation method, because the assessment process requires the clients to share their data with the server.
Another sensible approach to private evaluation in distributed settings is to query every client about her own evaluation score for the generative model, and then to aggregate the clients’ individual evaluation scores. While this aggregation-based approach needs every client to share only her score, the resulting ranking of generative models could be different from the centralized score-all evaluation. Therefore, a key question is how consistent a score aggregation-based evaluation could be with the score-all assessment.
In this work, we answer the mentioned question for the KID and FID evaluation metrics. In the KID case, we prove that accessing to the clients’ individual KID scores will provide sufficient information for ranking a group of generative models using the centralized KID-all metric. Specifically, we prove that the mean of clients’ KID scores, which we call KID-avg, will result in the same ranking of generative models as the potentially expensive-to-compute KID-all score. This result reveals the low communication and privacy costs of ranking the models according to the centralized KID-all score, which is identical to the ranking of the generative models using the averaged KID-avg score.
In contrast, we show that the knowledge of clients’ individual FID scores for a set of generative models is insufficient to compare those models according to the centralized FID-all score. We provide general examples of distributed scenarios in which the clients report the same FID values for two generative models, while the models lead to significantly different centralized FID scores. Furthermore, we show that the averaged FID scores across users, which we call the FID-avg score, would be minimized by a different generative model from the centralized-all minimizing model. Therefore, our theoretical results indicate the insufficiency of the clients’ individual FID scores for ranking the models according to their centralized FID-all score.
Finally, we discuss several numerical results on standard image datasets and generative model architectures which are consistent with our analytical results on the aggregation of the distance-based evaluation scores. We utilize the truncation technique to simulate a heterogeneous federated learning scenario where the clients’ image data have different distributions, and show several cases where two models share the same client-based FID scores, while their FID with respect to the collective set of clients’ data is significantly different. We also show the inconsistency between the FID-all and FID-avg’s rankings of a group of generative models. On the other hand, Our empirical results suggest that the centralized KID-all and distributed FID-avg aggregations provide consistent rankings of the tested generative models, while they could have different values. In the following, we summarize the main contributions of our study:
  • Studying the distance-based evaluation of generative models in heterogeneous distributed settings,
  • Proving the identical rankings given by the centralized KID-all score with the clients’ collective reference set and the KID-avg score averaging clients’ individual KIDs.
  • Demonstrating the insufficient knowledge of clients’ FID scores for ranking generative models based on the FID-all score with the clients’ collective reference dataset;
  • Presenting numerical results on the evaluation of generative models in distributed settings and empirically supporting the theoretical claims.
2 Related Work
A large body of related works [13] has focused on the evaluation of generative models in standard centralized learning settings. The existing evaluation scores can be categorized into two general groups: 1) Distance-based metrics defining a distance between the distribution of training data and learnt generative model. The distance between the real and fake distributions is usually computed after passing the samples through a pre-trained neural net offering a proper embedding of image data. The well-known evaluation scores in this category are the FID [4] and KID [5] scores. 2) Quality and diversity-based scores which output a score based on the sharpness and variety of the generated samples. The widely-used evaluation metrics in this category are the Inception score [14], precision and recall metrics [15, 16], and density and coverage scores [17]. We note that in this work our goal is not to introduce a new evaluation metric, and our aim is to analyze the extensions of these scores to distributed and federated learning problems under heterogeneous data distributions across clients.
In another set of related works, extensions of generative model training methods including GANs and diffusion models to distributed federated learning have been studied. [7] propose Fed-GAN to train GANs in a federated learning setting. In [9], the gradient from sample generation for the generator is exchanged on the server, while each client possesses a personalized discriminator. According to [10], different weights are assigned to local discriminators in the non-i.i.d. setting. Conversely, [11] ensure client privacy by sharing the discriminator across clients, while keeping the generator private. Additionally, [12] explore a dual diffusion paradigm to extend diffusion-based models into the federated learning setting, addressing concerns related to data leakage.
3 Preliminaries
In order to assess the performance of a generative model, a standard approach is to measure the distance between the distribution of real and generated data. Due to the high-dimensionality of standard image data, the evaluation of image-based generative models is typically performed after passing the data point through a pre-trained Inception model on the ImageNet dataset.
Specifically, a standard distance-based metric is the Fréchet inception distance (FID) defined as the $2$-Wasserstein distance between two Gaussian distributions with the mean and covariance parameters of the data distribution $P_{X}$, denoted by $\bm{\mu}_{X},C_{X}$, and with the mean and covariance of the generative model $P_{G}$, denoted by $\bm{\mu}_{G},C_{G}$:
$\mathrm{FID}(P_{X},P_{G})\>:=\>$ $\bigl{\|}\bm{\mu}_{X}-\bm{\mu}_{G}\bigr{\|}^{2}_{2}\>+\mathrm{Tr}\Bigl{(}C_{X}+C_{G}-2\bigl{(}C_{X}C_{G}\bigr{)}^{1/2}\Bigr{)}.$
Note that we can interpret the FID score as an approximation of the 2-Wasserstein distance given the first and second-order moments of the distributions.
Another widely-used distance-based score for the evaluation of generative models is the kernel inception distance (KID), which measures the maximum mean discrepancy (MMD) between the two distributions which is calculated using a kernel similarity function $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$. Here, the definition of the MMD distance between $P_{X}$ and $P_{G}$ based on kernel $k$ follows from
$\;\mathrm{KID}\bigl{(}P_{X},P_{G}\bigr{)}:=\;\sup_{\begin{subarray}{c}f\in\mathcal{H}_{k}:\,\|f\|_{\mathcal{H}_{k}}\leq 1\end{subarray}}\Bigl{(}\mathbb{E}_{X\sim P_{X}}\bigl{[}f(X)\bigr{]}-\mathbb{E}_{X^{\prime}\sim P_{G}}\bigl{[}f(X^{\prime})\bigr{]}\Bigr{)}^{2}$
In the above, $\mathcal{H}_{k}$ represents the reproducing kernel Hilbert space (RKHS) corresponding to kernel $k$, i.e., the function set that can be expressed by linear combination of kernel-based elements $k_{x}(\cdot)=k({x},\cdot)$ over $x\in\mathcal{X}$, and $\|\cdot\|_{\mathcal{H}_{k}}$ denotes the norm of this Hilbert space of functions.
4 Evaluation of Generative Models in Distributed Contexts
In this section, we discuss two extensions of distance-based evaluation scores from a centralized case to heterogeneous distributed learning settings. In our analysis, we use $\mathcal{D}(P_{X},P_{G})$ to denote a general distance between data distribution $P_{X}$ and the generative model $P_{G}$. Considering the standard distance-based metrics, $\mathcal{D}$ can be chosen to be the FID or KID score, which we will analyze in this section.
In a standard centralized setting, we have only a single distribution $P_{X}$ for real data. However, the main characteristic of a heterogeneous distributed learning problem is the multiplicity of the involved clients’ distribution. Here, we suppose a distributed setting with $k$ clients and use $P_{X_{1}},\ldots,P_{X_{k}}$ to denote their underlying distributions, i.e. $P_{X_{i}}$ stands for the data distribution at client $i$. In addition, we assume that every client $i$ has a fraction $0\leq\lambda_{i}\leq 1$ of the data in the network, i.e, $\lambda_{i}=\frac{n_{i}}{n}$ where $n=\sum_{j=1}^{k}n_{j}$ is the total number of samples in the network and $n_{i}$ is the number of samples at client $i$.
As a result of multiple input distributions, we need to define an aggregate evaluation score that is based on distance measure $\mathcal{D}$. The aggregate distance is supposed to summarize the performance of the generative model $P_{G}$ in only one score. To do this, we consider and analyze two reasonable ways of defining the aggregate score:
  1. 1.
    Collective-data-based Score $\mathcal{D}_{\mathrm{all}}$: The score-all with respect to the collective data of the clients is the distance between $P_{G}$ to the averaged distribution $\widehat{P}_{X}:=\sum_{i=1}^{k}\lambda_{i}P_{X_{i}}$:
    $\mathcal{D}_{\mathrm{all}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}\,:=\,\mathcal{D}\Bigl{(}\widehat{P}_{X}\,,\,P_{G}\Bigr{)}.$ (1)
  2. 2.
    Average Score $\mathcal{D}_{\mathrm{avg}}$: The score-avg is the mean of the client’s individual distance measures, i.e.
    $\mathcal{D}_{\mathrm{avg}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}\,:=\,\sum_{i=1}^{k}\lambda_{i}\mathcal{D}\bigl{(}P_{X_{i}},P_{G}\bigr{)}.$ (2)
In the above, note that $\sum_{i=1}^{k}\lambda_{i}P_{X_{i}}$ is indeed a mixture distribution with $k$ components $P_{X_{1}},\ldots,P_{X_{k}}$ with frequency parameters $\lambda_{1},\ldots,\lambda_{k}$. As discussed in the introduction, the score-all metric is the desired evaluation score following the literature on distributed generative modeling. However, due to the privacy and communication costs, the score-avg metric is cheaper to compute over the network. Therefore, understanding the connections between the rankings of the models following the two aggregate scores is the key to apply the score-avg as a surrogate for the score-all metric in the assessment process.
4.1 KID-based Evaluation in Distributed Settings
Here, we focus on the KID score and analyze the consistency of model rankings following KID-all and KID-avg scores. The following theorem proves that the KID-all and KID-avg are guaranteed to provide a consistent ordering of a group of generative models, and there is a monotonic relationship between the two scores independent of the choice of the generative model. We defer the theoretical statements’ proof to the Appendix.
Theorem 1.
Consider a kernel function $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ and the resulting KID score. Then for the clients’ distributions $P_{X_{1}},\ldots,P_{X_{k}}$ with frequency parameters $\lambda_{1},\ldots,\lambda_{k}$, we will have the following for the average distribution $\widehat{P}_{X}=\sum_{j=1}^{k}\lambda_{j}P_{X_{j}}$:
$\mathrm{KID}_{\mathrm{avg}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}\,=\,\mathrm{KID}_{\mathrm{all}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}+\sum_{i=1}^{k}\lambda_{i}\,\mathrm{KID}\bigl{(}\widehat{P}_{X},P_{X_{i}}\bigr{)},$
which implies a monotonic relationship between KID-all and KID-avg as a function of $P_{G}$.
The above theorem shows the consistent rankings implied by the aggregate KID-avg and KID-all scores, as the difference between the scores remains constant while changing the model $P_{G}$. Therefore, to rank a set of models based on the desired KID-all score, the clients can only share their own evaluated KID, and the server can average the reported scores to recover the models’ ranking according to the centralized KID-all.
4.2 FID-based Evaluation in Distributed Settings
In the case of FID score, we utilize the formulation of FID as the 2-Wassesrtein-distance between the Gaussian-fitted model that leads to a Riemannian geometry. This observation results in the following theorem on FID-all and FID-avg aggregations.
Theorem 2.
Suppose that $P_{X_{1}},\ldots,P_{X_{k}}$ are the clients’ distributions with the mean parameters $\bm{\mu}_{1},\ldots,\bm{\mu}_{k}$ and covariance matrices $C_{1},\ldots,C_{k}$, respectively, in the Inception net embedding. Then, the following holds for a generative model $P_{G}$ with mean $\bm{\mu}_{G}$ and covariance $C_{G}$.
  1. 1.
    For FID-all, if we define random $\widehat{X}$ with mean $\widehat{\bm{\mu}}=\sum_{i=1}^{k}\lambda_{i}\bm{\mu}_{i}$ and covariance matrix $\widehat{C}=\sum_{i=1}^{k}\lambda_{i}\bigl{(}C_{i}+\bm{\mu}_{i}\bm{\mu}_{i}^{\top}-\widehat{\bm{\mu}}\widehat{\bm{\mu}}^{\top}\bigr{)}$, then
    $\mathrm{FID}_{\mathrm{all}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}\,=\,\mathrm{FID}\bigl{(}P_{\widehat{X}},P_{G}\bigr{)}.$ (3)
  2. 2.
    For FID-avg, if we define $\widetilde{X}$ with mean $\widehat{\bm{\mu}}=\sum_{i=1}^{k}\lambda_{i}\bm{\mu}_{i}$ and covariance matrix $\widetilde{C}$ as the unique solution to $\widetilde{C}=\sum_{i=1}^{k}\lambda_{i}\bigl{(}\widetilde{C}^{1/2}C_{i}\widetilde{C}^{1/2}\bigr{)}^{1/2}$, then
    $\mathrm{FID}_{\mathrm{avg}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}\,=\,\mathrm{FID}\bigl{(}P_{\widetilde{X}},P_{G}\bigr{)}+\sum_{i=1}^{k}\lambda_{i}\mathrm{FID}\bigl{(}P_{\widetilde{X}},P_{X_{i}}\bigr{)}.$
Remark 1.
In Theorem 2, $\widetilde{C}$ represents the covariance of the Wasserstein barycenter of the Gaussian distributions. If $C_{1},\ldots,C_{k}$ commute, i.e. $\forall i,j:\>C_{i}C_{j}=C_{j}C_{i}$, it simplifies to $\widetilde{C}=\bigl{(}\sum_{i=1}^{k}\lambda_{i}C_{i}^{{1}/{2}}\bigr{)}^{2}$
In the above theorem, we show the optimal mean vectors for FID-all and FID-avg aggregations are the same. In contrast, the optimal covariance matrix of FID-avg denoted by $\widetilde{C}$ has no dependence on the choice of $\bm{\mu}_{i}$’s, while the optimal covariance matrix of FID-all $\widetilde{C}$ will be affected by the difference between $\bm{\mu}_{i}$’s due to the term $\sum_{i=1}^{k}\lambda_{i}(\bm{\mu}_{i}\bm{\mu}_{i}^{\top}-\widehat{\bm{\mu}}\widehat{\bm{\mu}}^{\top})$. Therefore, Theorem 2 shows the distance-minimizing covariance matrices for FID-all and FID-avg are different in heterogeneous settings where $\bm{\mu}_{i}$’s are not identical, which shows the existence of models inconsistently ranked by FID-all and FID-avg.
As the FID-avg aggregation may not provide the FID-all’s ranking, a natural question is whether there exists another aggregation of the individual $\mathrm{FID}(P_{X_{i}},P_{G})$ scores which maintains the ranking of centralized FID-all evaluation. The following proposition shows that the answer to this question is no and it is possible that a group of clients assign identical individual FID scores to two generative models $P_{G}$ and $P_{G^{\prime}}$, while $P_{G}$ and $P_{G^{\prime}}$ have different FID-all scores.
Proposition 1.
In Theorem 2’s setting, suppose there exists $i,j$ that $\bm{\mu}_{i}\neq\bm{\mu}_{j}$. Also, assume the number of clients is less than the dimension of Inception embedding. Then, if we consider the FID-all minimizing $\widehat{G}$ with $\widehat{\bm{\mu}}$ and $\widehat{C}$ parameters, there exists a different generator $G^{\prime}$ that shares the same client-assigned FID scores $\mathrm{FID}(P_{X_{i}},P_{G^{\prime}})=\mathrm{FID}(P_{X_{i}},P_{\widehat{G}})$ for all $i$, while
[…]
As a result, two generative models may receive the same FID scores from a group of clients, while their FID-all scores are different according to the same set of clients. Therefore, neither averaging nor any other aggregation of clients’ FID scores can always preserve the ranking of the FID-all metric. Note that this statement does not apply to the KID score as shown in Theorem 1.
Table 1: The individual FID-scores of each clients and two aggregated FID scores.
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 FID-avg FID-all
Generator 1 281.76 198.54 212.00 225.21 129.14 123.59 195.04 100.49
Generator 2 282.44 199.24 212.36 226.21 129.34 122.61 195.37 190.93
Difference 0.68 0.70 0.36 1.00 0.20 0.98 0.33 90.44
Figure 2: Visual examples of generators and clients in Table 1.
5 Numerical Results
In this section, we test the theoretical results discussed in the text on the distributed ranking of generative models using the FID and KID metrics. We apply the truncation technique [16] to simulate a distributed evaluation setting with heterogeneous data distributions. Utilizing the simulated distributed scenario, we conducted numerical experiments to assess our theoretical results on FID and KID-based distributed evaluation. Also, we examined an implication of our results on detecting mode collapse in a distributed setting using the FID-avg and KID-avg scores.
5.1 Distributed vs. Centralized Evaluation using the FID metric
We earlier discussed that under identical FID evaluation scores assigned by the clients to two models, the model’s centralized FID score (FID-all) could still take different values. This subsection details experiments conducted on the FFHQ dataset [18] to demonstrate the possibility of such an outcome in real distributed settings.
(a) Generator 1 ($\tau=0.01$)
(b) Generator 2 ($\tau=0.2$)
(c) Generator 3 ($\tau=0.4$)
(d) Generator 4 ($\tau=0.6$)
(e) Generator 5 ($\tau=0.8$)
(f) Generator 6 ($\tau=1.0$)
Figure 3: Illustration of generators with different truncation factors. $\tau$ is the truncation parameter.
(a) Client 1
(b) Client 2
(c) Client 3
(d) Client 4
(e) Client 5
(f) Client 6
Figure 4: Illustration of simulated clients with heterogeneous distributions via truncation technique.
Experiment Setup. We utilized the pre-trained StyleGAN2-ADA [18] generator on FFHQ [18] dataset provided in the official GitHub repository. The model’s generated images were of size $1024\times 1024$. Regarding the generators, by applying the truncation technique [16] and setting the truncation factor $\tau$ to vary over {0.01, 0.6}, we obtained two generators with highly different diversity scores; a higher truncation factor results in greater diversity. To simulate clients with highly heterogeneous distributions, we again applied the truncation technique and generated image samples centered around the randomly selected mean vectors. While the generated images for each simulated client looked similar, the samples of one client could be completely different from those of other simulated clients. We show randomly-selected samples from the generators and clients in Figure 2. Such a heterogeneous distributed scenario could occur in an actual mobile network, where the users’ collected photos have a higher inter-client variance than intra-client diversity. We calculated the individual FID score for each generator-client pair, averaging these to determine the FID-avg as in Equation 2. All the data in the users’ dataset were then used to calculate the FID-all metric for each generator as in Equation 1.
Experimental Results. As shown in Table 1, the experiment led to the identification of six clients with negligible differences in client-based FID scores across all clients (FID scores’ difference bounded by 1.0), suggesting that the two generators have similar FID-based scores from the perspective of each of the six clients. Therefore, both generators also had similar FID-avg scores (195.04 vs. 195.37). However, the FID-all scores (based on the six clients) led to significantly different scores for the two generators (100.49 vs. 190.93). The numerical result confirms our theoretical statement that similar client-based FID scores does not guarantee similar centralized FID-all scores.
(a) FID-scores
(b) KID-scores
(c) Ranking(FIDs)
(d) Ranking(FID×$\times$×KID)
Figure 5: The results of evaluating generators in Figure 3 over clients in Figure 4.
Figure 6: Mode collapse Detection in the distributed evaluation setting. FID-avg failed to detect mode collapse, while KID-avg succeeded.
5.2 Comparison between Rankings implied by FID-based and KID-based Distributed Scores
Our theoretical analysis shows that the distributed KID-avg changes monotonically with the centralized KID-all score. To validate this claim, we conducted experiments on the simulated FFHQ data [18]. We also tested the hypothesis on several other datasets including Gaussian mixture clients, AFHQ [18], CIFAR [19], and ImageNet [20, 21] discussed in the Appendix.
Experiment Setting. Following a similar methodology as described in the previous subsection, we synthesized a series of diversity-controlled generators by applying the standard truncation technique to the random noise vector $\mathbf{z}$ [16]. We varied the truncation factor $\tau$ over $[0.01,1.0]$. The effect of changing the truncation factor on the generated samples’ diversity is illustrated in Figure 3. For every attempted $\tau$, we generated 5k samples. Additionally, to simulate a distributed setting with heterogeneous data distributions, we simulated 100 clients, each with 100 images synthesized with truncation factor $\tau=0.25$. We evaluated the generators in the distributed setting of the simulated clients using both the discussed FID and KID-based evaluation scores.
Numerical Results. As shown in Figure 5a and Figure 5b, FID-avg and KID-avg led to different rankings of the models. The plot of FID-avg versus FID-all led to a U-shaped curve, revealing inconsistent rankings of the models. On the other hand, the difference between KID-avg and KID-all remained constant for the simulated generators, as shown in Theorem 1. These findings are also visible in the comparative rankings based on FID and KID scores in Figures 5c and 5d. The results suggest that although KID scores and FID-all perfer generators with higher truncation factors and diversity, FID-avg preferred generated data with limited diversity matching the bounded diversity level at each client.
Knowing that the averaged scores are less costly for computation in distributed settings, KID-avg seems to offer a reliable proxy of KID-all and FID-all for a cost-efficient distributed ranking of generative models.
5.3 Mode Collapse Detection in Distributed Settings via KID-avg
During the training of generative adversarial networks (GANs), mode collapse may happen when the model fails to capture the diversity in the true distribution[22]. We analyzed the behavior of KID-avg and FID-avg for detecting mode collapse. In our experiments, the target generative model is evaluated by distributed FID-avg and KID-avg scores at fixed intervals during training, utilizing the same 100 clients described in Section 5.2. We simulated an experimental setting that led to mode collapse where we changed the training data at an epoch and begin training over a single data point. We report the evaluation outcomes and provide examples of generated images in Figure 6. We also show the centralized evaluation score FID-all and KID-all as a reference.
As demonstrated in Figure 6, the evaluated FID-avg scores could not help with detecting mode collapse, whereas FID-all, KID-all, and KID-avg all revealed a sudden deterioration in the scores. As FID-all and KID-all scores are more expensive for computation in a distributed setting, KID-avg could provide a less costly metric to detect mode collapse.
6 Conclusion
In this paper, we studied the evaluation of generative models in distributed settings where the clients can have heterogeneous data distributions. We discussed the score aggregation for widely-used FID and KID metrics to find the ranking of a group of generators according to the centralized FID-all and KID-all evaluation. We proved that averaging clients’ KID scores will address this task for the KID-based evaluation. On the other hand, we showed that no aggregations of clients’ FID scores can always recover the ranking of a centralized FID evaluation of generative models. A potential future direction is to characterize the computational and communication costs of distributed ranking of generative models based on the centralized FID-all score. In addition, extending the study to other evaluation metrics in the literature will be an interesting topic for future study.
Limitations and Broader Impact
Our numerical experiments were limited to image-based generative models, and their extension to other domains such as video and text will be insightful on the application of KID-avg score for an inexpensive distributed evaluation of generative models.
Appendix A Proofs
A.1 Proof of Theorem 1
  1. 1.
    Note that according to the definition,
    $\mathrm{FID}_{\mathrm{all}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}\,=\,\mathrm{FID}\bigl{(}\sum_{i=1}^{k}\lambda_{i}P_{X_{i}},P_{G}\bigr{)}.$
    Since the FID score depends only on the the mean and covariance parameters in the Inception-based semantic space, we can replace $\sum_{i=1}^{k}\lambda_{i}P_{X_{i}}$ with any other distribution that shares the same mean and covariance parameters, and the FID value will not change. Observe that given mean parameters $\bm{\mu}_{1},\ldots,\bm{\mu}_{k}$, the Inception-based mean of $\sum_{i=1}^{k}\lambda_{i}P_{X_{i}}$ will be $\widehat{\bm{\mu}}=\sum_{i=1}^{k}\lambda_{i}\bm{\mu}_{i}$. Therefore, the Inception-based covariance matrix of $\sum_{i=1}^{k}\lambda_{i}P_{X_{i}}$ follows from
    $\sum_{i=1}^{k}\lambda_{i}\mathbb{E}_{P_{i}}\bigl{[}\bigl{(}X_{i}-\widehat{\bm{\mu}}\bigr{)}\bigl{(}X_{i}-\widehat{\bm{\mu}}\bigr{)}^{\top}\bigr{]}\>$ $=\>\sum_{i=1}^{k}\lambda_{i}\Bigl{[}C_{i}+\bigl{(}\bm{\mu}_{i}-\widehat{\bm{\mu}}\bigr{)}\bigl{(}\bm{\mu}_{i}-\widehat{\bm{\mu}}\bigr{)}^{\top}\Bigr{]}$
    $=\>\sum_{i=1}^{k}\lambda_{i}\Bigl{[}C_{i}+\bm{\mu}_{i}\bm{\mu}_{i}^{\top}\Bigr{]}-\widehat{\bm{\mu}}\widehat{\bm{\mu}}^{\top}$
    $=\>\widehat{C}.$
    Therefore, since we assume $\widehat{X}$ has the Inception-based mean and covariance $\widehat{\bm{\mu}}$ and $\widehat{C}$, the proof of this part is complete.
  2. 2.
    According to the definition, FID-avg can be written as
    $\mathrm{FID}_{\mathrm{avg}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}\,\>:=\>\sum_{i=1}^{k}\lambda_{i}\mathrm{FID}\bigl{(}P_{X_{i}},P_{G}\bigr{)}.$
    Therefore, we have
    $\mathrm{FID}_{\mathrm{avg}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}\,$
    =$\stackrel{{\scriptstyle(a)}}{{=}}\>$ $\sum_{i=1}^{k}\lambda_{i}W^{2}_{2}\Bigl{(}\mathcal{N}(\bm{\mu}_{i},C_{i}),\mathcal{N}(\bm{\mu}_{G},C_{G})\Bigr{)}$
    =$\stackrel{{\scriptstyle(b)}}{{=}}\>$ $\sum_{i=1}^{k}\lambda_{i}\Bigl{[}\|\bm{\mu}_{i}-\bm{\mu}_{G}\|^{2}_{2}+\mathrm{Tr}\bigl{(}C_{i}+C_{G}-(C_{i}C_{G})^{1/2}\bigr{)}\Bigr{]}$
    =\>= $\sum_{i=1}^{k}\Bigl{[}\lambda_{i}\|\bm{\mu}_{i}-\bm{\mu}_{G}\|^{2}_{2}\Bigr{]}+\sum_{i=1}^{k}\Bigl{[}\lambda_{i}\mathrm{Tr}\bigl{(}C_{i}+C_{G}-(C_{i}C_{G})^{1/2}\bigr{)}\Bigr{]}$
    =$\stackrel{{\scriptstyle(c)}}{{=}}\>$ $\|\widehat{\bm{\mu}}-\bm{\mu}_{G}\|^{2}_{2}+\sum_{i=1}^{k}\Bigl{[}\lambda_{i}\|\widehat{\bm{\mu}}-\bm{\mu}_{i}\|^{2}_{2}\Bigr{]}$
    $\qquad+\mathrm{Tr}\bigl{(}C_{G}+\widehat{C}-(C_{G}\widehat{C})^{1/2}\bigr{)}+\sum_{i=1}^{k}\Bigl{[}\lambda_{i}\mathrm{Tr}\bigl{(}C_{i}+\widehat{C}-(C_{i}\widehat{C})^{1/2}\bigr{)}\Bigr{]}$
    =\>= $\|\widehat{\bm{\mu}}-\bm{\mu}_{G}\|^{2}_{2}+\mathrm{Tr}\bigl{(}C_{G}+\widehat{C}-(C_{G}\widehat{C})^{1/2}\bigr{)}$
    $\qquad+\sum_{i=1}^{k}\Bigl{[}\lambda_{i}\|\widehat{\bm{\mu}}-\bm{\mu}_{i}\|^{2}_{2}+\lambda_{i}\mathrm{Tr}\bigl{(}C_{i}+\widehat{C}-(C_{i}\widehat{C})^{1/2}\bigr{)}\Bigr{]}$
    =\>= $\|\widehat{\bm{\mu}}-\bm{\mu}_{G}\|^{2}_{2}+\mathrm{Tr}\bigl{(}C_{G}+\widehat{C}-(C_{G}\widehat{C})^{1/2}\bigr{)}$
    $\qquad+\sum_{i=1}^{k}\lambda_{i}\Bigl{[}\|\widehat{\bm{\mu}}-\bm{\mu}_{i}\|^{2}_{2}+\mathrm{Tr}\bigl{(}C_{i}+\widehat{C}-(C_{i}\widehat{C})^{1/2}\bigr{)}\Bigr{]}$
    =$\stackrel{{\scriptstyle(d)}}{{=}}\>$ $\mathrm{FID}(P_{\widehat{X}},P_{G})+\sum_{i=1}^{k}\lambda_{i}\mathrm{FID}(P_{\widehat{X}},P_{X_{i}}).$
    In the above, $(a)$ follows from the Wasserstein-based definition of FID distance. $(b)$ comes from the well-known closed-form expression of the 2-Wasserstein distance between Gaussian distributions [23]. $(c)$ is the result of applying the weighted barycenter of vector $\bm{\mu}_{1},\ldots,\bm{\mu}_{k}$ that can be seen to be $\widehat{\bm{\mu}}$ and the weighted barycenter of positive semi-definite covariance matrices $C_{1},\ldots,C_{k}$ that has been shown to be the unique matrix $\widehat{C}$ that solves the equation $\widetilde{C}=\sum_{i=1}^{k}\lambda_{i}\bigl{(}\widetilde{C}^{1/2}C_{i}\widetilde{C}^{1/2}\bigr{)}^{1/2}$ [24, 25]. $(d)$ is the direct consequence of the Wasserstein-based definition of the FID distance and the closed-form expression of the 2-Wasserstein distance between Gaussians. Therefore, the proof is complete.
A.2 Proof of Theorem 2
To show this theorem, we note that if $\phi(X)$ is the kernel feature map for kernel $k$ used to define the KID distance, i.e. $k(x,y)=\langle\phi(x),\phi(y)\rangle$ is the inner product of the feature maps applied to $x,y$, then it can be seen that the kernel-$k$-based MMD distance can be written as
$\mathrm{MMD}\bigl{(}P_{X},P_{G}\bigr{)}:=$ $\mathbb{E}_{X,X^{\prime}\sim P_{X}}\bigl{[}k(X,X^{\prime})\bigr{]}+\mathbb{E}_{Y,Y^{\prime}\sim P_{G}}\bigl{[}k(Y,Y^{\prime})\bigr{]}-2\,\mathbb{E}_{X\sim P_{X},\,Y\sim P_{G}}\bigl{[}k(X,Y)\bigr{]}$
$=$ $\Bigl{\|}\mathbb{E}\bigl{[}\phi(X)\bigr{]}-\mathbb{E}\bigl{[}\phi(Y)\bigr{]}\Bigr{\|}^{2}.$
Therefore, following the definition of KID-avg, we can write
$\mathrm{KID}_{\mathrm{avg}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}\,\>:=$ $\>\sum_{i=1}^{k}\lambda_{i}\mathrm{KID}\bigl{(}P_{X_{i}},P_{G}\bigr{)}$
$=$ $\>\sum_{i=1}^{k}\lambda_{i}\mathrm{MMD}_{\phi}\bigl{(}P_{X_{i}},P_{G}\bigr{)}$
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $\>\sum_{i=1}^{k}\lambda_{i}\Bigl{\|}\mathbb{E}\bigl{[}\phi(X_{i})\bigr{]}-\mathbb{E}\bigl{[}\phi(G(Z))\bigr{]}\Bigr{\|}^{2}$
=$\stackrel{{\scriptstyle(b)}}{{=}}$ $\>\bigl{\|}\mathbb{E}\bigl{[}\phi(\widehat{X})\bigr{]}-\mathbb{E}\bigl{[}\phi(G(Z))\bigr{]}\bigl{\|}^{2}+\sum_{i=1}^{k}\Bigl{[}\lambda_{i}\bigl{\|}\mathbb{E}\bigl{[}\phi(X_{i})\bigr{]}-\mathbb{E}\bigl{[}\phi(\widehat{X})\bigr{]}\bigr{\|}^{2}\Bigr{]}$
=$\stackrel{{\scriptstyle(c)}}{{=}}$ $\>\mathrm{MMD}_{\phi}\bigl{(}\widehat{P}_{X},P_{G}\bigr{)}+\sum_{i=1}^{k}\Bigl{[}\lambda_{i}\mathrm{MMD}_{\phi}\bigl{(}\widehat{P}_{X},P_{X_{i}}\bigr{)}\Bigr{]}$
=$\stackrel{{\scriptstyle(d)}}{{=}}$ $\>\mathrm{KID}\bigl{(}\widehat{P}_{X},P_{G}\bigr{)}+\sum_{i=1}^{k}\Bigl{[}\lambda_{i}\mathrm{KID}\bigl{(}\widehat{P}_{X},P_{X_{i}}\bigr{)}\Bigr{]}$
=$\stackrel{{\scriptstyle(e)}}{{=}}$ $\>\mathrm{KID}_{\mathrm{all}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}+\sum_{i=1}^{k}\lambda_{i}\mathrm{KID}\bigl{(}\widehat{P}_{X},P_{X_{i}}\bigr{)}.$
[…]
A.3 Proof of Proposition 1
Consider the FID-all-minimizing parameters in Theorem 2 resulting in
$\mathrm{FID}_{\mathrm{all}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{\widehat{G}}\Bigr{)}\,$ $=\,\mathrm{FID}\bigl{(}\mathcal{N}(\widehat{\bm{\mu}},\widehat{C}),\mathcal{N}({\bm{\mu}_{\widehat{G}}},{C}_{\widehat{G}})\bigr{)}$
$=\,\bigl{\|}\widehat{\bm{\mu}}-\bm{\mu}_{\widehat{G}}\bigr{\|}^{2}_{2}+\mathrm{Tr}\Bigl{(}\widehat{C}+C_{\widehat{G}}-2\bigl{(}\widehat{C}C_{\widehat{G}}\bigr{)}^{1/2}\Bigr{)}.$
Note that since we assume the number of clients $k$ is less than the dimension of the embedding, there exists a unit-norm vector $\bm{\beta}$ ($\|\bm{\beta}\|_{2}=1$) in the embedding space that is orthogonal to all mean vectors $\bm{\mu}_{1},\ldots,\bm{\mu}_{k}$ and hence to their mean $\widehat{\bm{\mu}}=\frac{1}{k}\sum_{i=1}^{k}\bm{\mu}_{i}$. Given $u=\mathrm{Tr}\bigl{(}\sum_{i=1}^{k}\lambda_{i}\bigl{(}\bm{\mu}_{i}\bm{\mu}_{i}^{\top}-\widehat{\bm{\mu}}\widehat{\bm{\mu}}^{\top}\bigr{)}\bigr{)}$, we then consider the generator $G^{\prime}$ with the following mean and covariance parameters:
$\bm{\mu}_{G^{\prime}}\,=\,\widehat{\bm{\mu}}+\sqrt{u}\bm{\beta},\qquad C_{G^{\prime}}=C_{\widehat{G}}-\sum_{i=1}^{k}\lambda_{i}\bigl{(}\bm{\mu}_{i}\bm{\mu}_{i}^{\top}-\widehat{\bm{\mu}}\widehat{\bm{\mu}}^{\top}\bigl{)}\,=\,\sum_{i=1}\lambda_{i}C_{i}.$
We claim that the generators $\widehat{G}$ and $G^{\prime}$ lead to the same client-based FID scores as for every $i$
$\mathrm{FID}\bigl{(}P_{X_{i}},P_{G^{\prime}}\bigr{)}\,$ $=\,\mathrm{FID}\bigl{(}\mathcal{N}({\bm{\mu}}_{i},{C}_{i}),\mathcal{N}({\bm{\mu}_{{G}^{\prime}}},{C}_{{G}^{\prime}})\bigr{)}$
$=\,\bigl{\|}{\bm{\mu}}_{i}-\bm{\mu}_{{G}^{\prime}}\bigr{\|}^{2}_{2}+\mathrm{Tr}\Bigl{(}{C}_{i}+C_{{G}^{\prime}}-2\bigl{(}{C}_{i}C_{{G}^{\prime}}\bigr{)}^{1/2}\Bigr{)}$
$=\bigl{\|}{\bm{\mu}}_{i}-\bm{\mu}_{\widehat{G}}\bigr{\|}^{2}_{2}+u+\mathrm{Tr}\Bigl{(}{C}_{i}+C_{{G}^{\prime}}-2\bigl{(}{C}_{i}C_{\widehat{G}}\bigr{)}^{1/2}\Bigr{)}-u$
$=\bigl{\|}{\bm{\mu}}_{i}-\bm{\mu}_{\widehat{G}}\bigr{\|}^{2}_{2}+\mathrm{Tr}\Bigl{(}{C}_{i}+C_{{G}^{\prime}}-2\bigl{(}{C}_{i}C_{\widehat{G}}\bigr{)}^{1/2}\Bigr{)}$
$=\mathrm{FID}\bigl{(}P_{X_{i}},P_{\widehat{G}}\bigr{)}.$
On the other hand, for the FID-all of $G^{\prime}$ we have
$\mathrm{FID}_{\mathrm{all}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{{G}^{\prime}}\Bigr{)}\,$ $=\,\mathrm{FID}\bigl{(}\mathcal{N}(\widehat{\bm{\mu}},\widehat{C}),\mathcal{N}({\bm{\mu}_{{G}^{\prime}}},{C}_{{G}^{\prime}})\bigr{)}$
$=\,\bigl{\|}\widehat{\bm{\mu}}-\bm{\mu}_{{G}^{\prime}}\bigr{\|}^{2}_{2}+\mathrm{Tr}\Bigl{(}\widehat{C}+C_{{G}^{\prime}}-2\bigl{(}\widehat{C}C_{{G}^{\prime}}\bigr{)}^{1/2}\Bigr{)}$
$=\,\bigl{\|}\widehat{\bm{\mu}}-\bm{\mu}_{\widehat{G}}\bigr{\|}^{2}_{2}+u\mathrm{Tr}\Bigl{(}\widehat{C}+C_{\widehat{G}}-2\bigl{(}\widehat{C}C_{\widehat{G}}\bigr{)}^{1/2}\Bigr{)}+u$
$=\,\bigl{\|}\widehat{\bm{\mu}}-\bm{\mu}_{\widehat{G}}\bigr{\|}^{2}_{2}+\mathrm{Tr}\Bigl{(}\widehat{C}+C_{\widehat{G}}-2\bigl{(}\widehat{C}C_{\widehat{G}}\bigr{)}^{1/2}\Bigr{)}$
$\;\;+2\mathrm{Tr}\bigl{(}\sum_{i=1}^{k}\lambda_{i}\bigl{(}\bm{\mu}_{i}\bm{\mu}_{i}^{\top}-\widehat{\bm{\mu}}\widehat{\bm{\mu}}^{\top}\bigr{)}\bigr{)}$
$=\,\mathrm{FID}_{\mathrm{all}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{\widehat{G}}\Bigr{)}$
$\;\;+2\mathrm{Tr}\bigl{(}\sum_{i=1}^{k}\lambda_{i}\bigl{(}\bm{\mu}_{i}\bm{\mu}_{i}^{\top}-\widehat{\bm{\mu}}\widehat{\bm{\mu}}^{\top}\bigr{)}\bigr{)}$
Therefore, Proposition 1’s proof is complete.
Appendix B General 1-Wasserstein-Distance evaluation Metrics
Let $\mathbb{P}_{g}$ and $\mathbb{P}_{t}$ represent the distribution of generated set and training set. The Wasserstain-1 distance between $\mathbb{P}_{g}$ and $\mathbb{P}_{t}$ is,
$W(\mathbb{P}_{g},\mathbb{P}_{t})=\mathop{inf}\limits_{\lambda\in\Pi(\mathbb{P}_{g},\mathbb{P}_{t})}\mathbb{E}_{(x,y)\sim\lambda}[\|x-y\|],$ (4)
where $\Pi(\mathbb{P}_{g},\mathbb{P}_{t})$ denotes the set of all joint distribution $\lambda(x,y)$ whose marginal distribution are respectively $\mathbb{P}_{g}$ and $\mathbb{P}_{t}$. However, the direct estimation of $W(\mathbb{P}_{g},\mathbb{P}_{t})$ is highly intractable. On the other hand, the Kantorovich-Rubinstein duality [23] gives,
$W(\mathbb{P}_{g},\mathbb{P}_{t})=\mathop{sup}\limits_{\|f\|_{L}\leq 1}\mathbb{E}_{x\sim\mathbb{P}_{g}}[f(x)]-\mathbb{E}_{x\sim\mathbb{P}_{t}}[f(x)],$ (5)
where the supremum is over all the 1-Lipschitz functions $f:\mathcal{R}^{n}\rightarrow\mathcal{R}$. Therefore, if we have a parameterized family of functions $\{f_{\theta}\}_{\theta\in\Theta}$ that a 1-Lipschitz, we could considering solve this problem,
$\mathop{max}\limits_{\theta\in\Theta}\mathbb{E}_{x\sim\mathbb{P}_{g}}[f_{\theta}(x)]-\mathbb{E}_{x\sim\mathbb{P}_{t}}[f_{\theta}(x)].$ (6)
To estimate the supremum of Equation 5, we employ a family of non-linear neural network $f_{\theta}$ which are repeatedly stacked by the fully connected layer, the spectral normalization and RELU activation layer. There are three repeated blocks in the network $f_{\theta}$ and the last block does have RELU. The feature is extracted by pre-trained Inception-V3 network. By optimizing the parameters in $f_{\theta}$ to maximize $\mathbb{E}_{x\sim\mathbb{P}_{g}}[f_{\theta}(x)]-\mathbb{E}_{x\sim\mathbb{P}_{t}}[f_{\theta}(x)]$ over $\mathbb{P}_{g}$ and $\mathbb{P}_{t}$, we can finally get an estimation of $W(\mathbb{P}_{g},\mathbb{P}_{t})$. And similarly, we can also define average score W-avg and collective-data-based score W-all under the distributed learning setting. Similar to the CIFAR100 experiment in the main body of paper, we extracted samples from each single class of CIFAR100 and evaluate these samples on federated CIFAR10 dataset. We illustrate a subset of W-avg / W-all pairs in Figure 7. According to experiment results, we find that general 1-Wasserstein-Distance evaluation metric also shows inconsistent behaviours in the distributed evaluation settings.
Figure 7: Evaluation with 1-Wasserstein-Distance on Federated CIFAR10.
Appendix C Extra Experiment Results on FFHQ and AFHQ
C.1 Extra samples from clients
We show more samples from clients described in Section 5.2 on Figure 8.
(a) Client 7
(b) Client 8
(c) Client 9
(d) Client 10
(e) Client 11
(f) Client 12
Figure 8: Illustration of random samples from randomly selected variance-limited clients.
C.2 Results on AFHQ dataset
Utilizing a another model weight that is pre-trained on AFHQ-wild dataset [26], we extend the experiment described in Section 5.2. Similar to the main text, we gradually increase the truncation parameter of generators We show some examples from generators, clients on Figure 10 and Figure 9 and plot the relationship between FID-scores and KID-scores on Figure 11. The results that are illustrated in the figures still indicates that FID-avg can lead to different ranking with FID-all, while KID-avg can be treated as a more stable evaluation metric in the distributed learning setting.
Figure 9: Illustration of randomly-generated samples from the variance-controlled generators on AFHQ dataset. Images in each row are synthesized with the same truncation parameter.
(a) Client 1
(b) Client 2
(c) Client 3
(d) Client 4
(e) Client 5
(f) Client 6
Figure 10: Illustration of random samples from randomly selected variance-limited clients on AFHQ dataset.
Figure 11: Evaluation on AFHQ dataset with StyleGAN2.
Appendix D Evaluation on Synthetic Gaussian Mixture Datasets
Figure 12: Experimental results of Gaussian mixture dataset. (a): The optimal $\mathrm{var}_{x}$ parameters are different under FD-avg and FD-all evaluations. (b): Distance between KD-avg and KD-all remains the same.(c): the clients’ and generator’s samples.
As discussed in main text, the optimal selection of the covariance matrix differs for the FID-all and FID-avg aggregate scores. To illustrate this distinction, we performed a toy experiment, revealing that FID-avg attains its minimum value when the generator’s variance closely approximates that of an individual client, whereas FID-all is minimized when the variance equals that of the aggregate distribution.
Setup. Our experimental setup involves two clients, denoted by $C_{1}$ and $C_{2}$. $C_{1}$ possesses a dataset consisting of 50,000 samples drawn from the Gaussian distribution $\mathcal{N}([1,0]^{\top},\bm{\Sigma})$, while $C_{2}$ holds a dataset with 50,000 samples drawn from $\mathcal{N}([-1,0]^{\top},\bm{\Sigma})$, where $\bm{\Sigma}=\text{diag}([1,1]^{T})$. We introduce a generator, denoted as $G_{\mathrm{var}_{x}}$, which is parameterized by $\mathrm{var}_{x}$. $\mathrm{var}_{x}$ regulates the variance of the generator along the X-axis. Specifically, $G_{\mathrm{var}_{x}}$ generates 50,000 data points following a Gaussian distribution $\mathcal{N}([0,0]^{T},\bm{\Sigma_{G}})$, where $\bm{\Sigma_{G}}=\text{diag}([\mathrm{var}_{x},1]^{T})$. The relationship between the two clients and the generator is visually depicted in Figure 12. Additionally, we introduce an "ideal estimator" denoted as $\hat{E}=C_{1}\cup C_{2}$. This ideal estimator possesses the unique ability to replicate the distribution of the training dataset perfectly. We employ the ideal estimator as a reference for our analysis.
[…]
Results. By increasing $\mathrm{var}_{x}$ from 0 to 4, we get a sequence of FD-avg / FD-all pairs and we plot them with the $\mathrm{var}_{x}$ in Figure 12. Our experimental results highlight the following conclusions. First, we observed that the minimum of FD-all occurs at $\mathrm{var}_{x}=2$, while that of FD-avg occurs at $\mathrm{var}_{x}=1$, which indicates that the optimal solutions of $\mathrm{var}_{x}$ to minimize FD-all and FD-avg are inconsistent. In this case, FD-all and FD-avg lead to different rankings of the models with $\mathrm{var}_{x}=1$ and $\mathrm{var}_{x}=2$. Additionally, we observed that, counterintuitively, the ‘ideal estimator’ did not reach the minimum average of the Fréchet distances. The distance between KD-avg and KD-all remains the same with the change of $\mathrm{var}_{x}$ and both of which reach minimum at $\mathrm{var}_{x}=2$. The toy experiment highlights how a co-variance mismatch between clients and the collective dataset leads to inconsistent rankings according to aggregate Fréchet distances.
The Log-Likelihood Score. We also evaluated the synthetic Gaussian mixture dataset with the standard log-likelihood (LL) score. In this experiment, we note that we have access to the probability density functions (PDF) of the simulated generator. We utilized the generator $G_{var_{x}}$ described in the main text and performed the evaluation over the parameter $\mathrm{var}_{x}$ in the range $[0,40]$. As can be shown in the general case, LL-avg and LL-all led to the same value for every evaluated model. As shown in Figure 13a, they reached their maximum value at $\mathrm{var}_{x}=2$. On the other hand, we set a new generator $G_{mean_{x}}$ generating samples according to $\mathcal{N}([\mathrm{mean}_{x},0]^{\top},\bm{\Sigma})$, where $\bm{\Sigma}=\text{diag}([2,1]^{T})$. We gradually increased $\mathrm{mean}_{x}$ from -2 to 2 and plotted LL-avg, LL-all, and LL-ref in Figure 13b.
(a) $G_{var_{x}}$
(b) $G_{mean_{x}}$
Figure 13: Evaluation of synthetic Gaussian data with the aggregate log-likelihood scores.
Appendix E Evaluation on Different Federated Image Datasets
E.1 Experiment Settings
We evaluated our theoretical results on standard image datasets. In our experiments, we simulated heterogeneous federated learning experiments consisting of non-i.i.d. data at different clients: for CIFAR-10 [19], we considered 10 clients, each owning 5000 samples exclusively from a single class of the image dataset. Therefore, every client’s dataset contains 5000 images having the same label. Similar to the federated CIFAR10, federated CIFAR100 and federated ImageNet-32 are conducted by grouping samples from a each class.
Figure 14: Left: Images generated by a generative model obtain a lower FID-all. Right: Images from real datasets with class ’airplane’ obtain a lower FID-avg. FID-avg and FID-all lead to inconsistent rankings, while KID-avg and KID-all result in the same ranking.
Neural Net-based Generators. We have trained WGAN-GP [14] and DDPM [3] in a federated learning setting by utilizing FedAvg approach [6]. The experiment protocols for WGAN-GP and DDPM are copied from original works. The communication interval of FedAvg is set as 160 iterations for both WGAN-GP and DDPM. We have tried different communication intervals for both models. The communication frequency will affect model performance but have no influence on the conclusions in the main part of our paper.
Perfect Data-simulating Generators. In our CIFAR-10 experiments, we also simulated and evaluated an "ideal generator" capable of perfectly replicating all samples belonging to the ’airplane’ class in CIFAR10. In this scenario, the samples "generated" by the ideal generator exhibit impeccable fidelity but lack diversity since no samples from other categories can be produced.
E.2 FID-based and KID-based Evaluation of Generative Models.
We evaluated the generative models according to FID-all, FID-avg, KID-all, and KID-avg as defined in Section 4. In several cases, we observed that FID-all / FID-avg could assign inconsistent rankings to the generators. Specifically, we computed FID-all and FID-avg for the ideal ’airplane’-class-based generator and neural net-based DDPM generators under the distributed CIFAR10 setting. We present some examples generated from the two generators in Figure 14 and report their scores according to the four metrics. The results suggest that FID-avg assigns a considerably lower score to the ideal ’airplane’-based generator, whose images preserve perfect details but lack diversity in image categories. Conversely, FID-all assigns a relatively lower value to the DDPM model because its images possess greater diversity. On the other hand, we also observed that KID-avg and KID-all give consistent rankings. Both of them led to the evaluation that the ideal plane generator is slightly better than the DDPM generator. In our implementation of KID-scores, we utilized the standard implementation of KID measurement from data with a polynomial kernel, $k(\bf{x},\bf{y})=\left(\frac{1}{d}\bf{x}^{T}\bf{y}+1\right)^{3}$, where $d$ is the dimension of feature vector. We note that our theoretical finding on the evaluation consistency under KID-all and KID-avg applies to every kernel similarity function. We also test KID-scores with a Gaussian RBF kernel $k_{\sigma}^{\text{rbf}}(\bf{x},\bf{y})=\exp\left(-\frac{1}{2\sigma^{2}}\|\bf{x}-\bf{y}\|^{2}\right)$ as formulated in [5], where we chose $\sigma=\sqrt{d}$ in the experiments. For images generated by diffusion model $\text{KID}^{\text{rbf}}$-all gives $4.277e^{-3}$ while $\text{KID}^{\text{rbf}}$-avg gives $4.295e^{-3}$. And for the airplane images in CIFAR10, $\text{KID}^{\text{rbf}}$-all gives $4.283e^{-3}$ while $\text{KID}^{\text{rbf}}$-avg gives $4.301e^{-3}$. The results indicate that for Gaussian RBF kernel $k_{\sigma}^{\text{rbf}}$, $\text{KID}^{\text{rbf}}$-all and $\text{KID}^{\text{rbf}}$-avg still gives consistent results. In this case, the $\text{KID}^{\text{rbf}}$-based evaluation suggests the images sampled from the diffusion model have higher quality than the set of airplane images in the CIFAR10 dataset.
E.3 Evaluate Sequence of Net-based Generator on Federated CIFAR10
We trained the WGAN-GP[27] generative models multiple times using different random states, and we set different training lengths for every training procedure. We saved the models at different checkpoints every $10$ epochs, which is common in training generative models to select the best-performing saved model according to an evaluation metric.
Our numerical results suggest that the gap between KID-all and KID-avg remains constant and hence they lead to the same rankings of the generative models. Here, we conducted our evaluations on all the generative models instances as previously described, and the results are visualized in the left sub-figure of Figure 15. These findings reveal that all distinct generators consistently exhibit a uniform gap between KID-avg and KID-all. Consequently, our results indicate that the rankings established by KID-avg consistently align with those of KID-all in distributed learning settings.
Figure 15: Left: KID evaluations of WGAN-GP checkpoints on federated CIFAR10. Right: FID-based evaluations on federated CIFAR-10.
E.4 Evaluate CIFAR100 Generator on Federated CIFAR10
To further experiment the ranking of generative models according to the discussed aggregate scores, we extracted samples from each class of CIFAR-100 and treated them as the output of one hundred distinct generators, each corresponding to a single class. By assessing these generators on the federated CIFAR-10 dataset, we obtained one hundred pairs of FID-avg / FID-all values, and a subset of these pairs with inconsistent rankings according to FID-all/FID-avg is visualized in the right of Figure 15. The complete set of evaluation results is available on Table 2. These results further highlight that the rankings provided by FID-all and FID-avg can exhibit inconsistencies in the context of distributed learning. Such inconsistencies could pose a challenge when selecting from a series of checkpoints or model architectures during the training of generative models in distributed learning scenarios, where a distributed computation of FID-all is more challenging than obtaining FID-avg due to privacy considerations.
Class FID-all FID-avg KID-all KID-avg Class FID-all FID-avg KID-all KID-avg
0 267.4 285.9 0.201 0.253 25 142.4 173.2 0.067 0.116
1 173.2 205.5 0.109 0.165 26 139.4 175.5 0.063 0.114
2 151.8 185.4 0.072 0.123 27 114.5 153.7 0.054 0.102
3 124.3 162.2 0.047 0.100 28 182.0 205.0 0.112 0.160
4 117.0 156.1 0.049 0.100 29 143.3 185.2 0.083 0.133
5 157.6 185.2 0.088 0.138 30 147.4 179.8 0.084 0.134
6 142.1 179.7 0.062 0.115 31 160.4 193.6 0.108 0.158
7 144.9 179.2 0.066 0.113 32 115.4 156.4 0.032 0.085
8 155.0 187.7 0.085 0.135 33 146.9 181.2 0.077 0.126
9 180.7 203.6 0.106 0.156 34 127.5 163.4 0.068 0.117
10 183.2 208.8 0.099 0.151 35 150.4 184.7 0.080 0.131
11 151.9 184.9 0.073 0.127 36 142.5 176.5 0.070 0.126
12 126.1 163.1 0.056 0.108 37 126.7 162.9 0.066 0.118
13 127.0 159.4 0.058 0.112 38 110.5 151.3 0.050 0.100
14 145.8 182.9 0.071 0.126 39 233.5 257.2 0.145 0.196
15 124.2 165.2 0.054 0.107 40 158.9 187.6 0.078 0.128
16 190.3 213.6 0.112 0.161 41 151.1 181.0 0.072 0.126
17 161.5 192.8 0.117 0.165 42 137.2 174.7 0.071 0.122
18 142.7 179.8 0.069 0.118 43 152.0 186.0 0.090 0.139
19 112.4 154.1 0.047 0.101 44 126.7 165.8 0.050 0.103
20 194.2 215.9 0.123 0.175 45 135.4 173.0 0.059 0.111
21 164.8 194.9 0.099 0.148 46 155.1 187.1 0.080 0.133
22 205.1 227.0 0.126 0.180 47 174.2 203.4 0.118 0.169
23 191.1 219.7 0.130 0.186 48 145.0 175.3 0.077 0.129
24 174.9 202.8 0.105 0.155 49 164.8 194.6 0.110 0.164
50 113.8 154.0 0.045 0.095 75 151.7 184.3 0.096 0.145
51 150.4 184.6 0.073 0.125 76 151.5 183.5 0.068 0.118
52 195.3 222.1 0.167 0.218 77 139.0 174.9 0.066 0.117
53 279.7 299.1 0.217 0.270 78 197.3 228.3 0.130 0.182
54 170.2 201.3 0.098 0.149 79 138.9 174.8 0.062 0.114
55 104.7 146.3 0.034 0.086 80 111.1 149.9 0.043 0.095
56 144.2 178.5 0.075 0.126 81 131.7 165.3 0.068 0.119
57 192.5 219.9 0.115 0.163 82 197.6 227.2 0.123 0.178
58 131.5 161.0 0.067 0.121 83 202.6 230.6 0.122 0.173
59 149.7 183.7 0.093 0.144 84 144.5 177.3 0.065 0.111
60 188.0 216.0 0.144 0.197 85 123.0 160.7 0.079 0.125
61 249.2 270.6 0.179 0.229 86 168.0 193.0 0.083 0.135
62 202.1 230.6 0.133 0.184 87 170.5 196.0 0.098 0.152
63 140.6 175.3 0.071 0.120 88 133.4 170.6 0.068 0.120
64 118.7 156.1 0.049 0.100 89 122.3 158.1 0.059 0.112
65 102.2 142.2 0.024 0.077 90 110.7 148.3 0.041 0.093
66 121.7 159.1 0.054 0.105 91 124.0 160.7 0.048 0.098
67 132.5 167.8 0.063 0.115 92 175.4 206.2 0.096 0.149
68 139.7 173.1 0.073 0.123 93 129.7 166.3 0.054 0.109
69 143.2 176.0 0.068 0.121 94 213.4 235.2 0.162 0.212
70 178.4 209.5 0.095 0.148 95 154.7 185.2 0.084 0.133
71 169.4 199.0 0.120 0.167 96 147.8 181.7 0.090 0.138
72 114.1 155.6 0.043 0.094 97 137.1 171.4 0.068 0.119
73 137.9 170.8 0.071 0.124 98 157.1 188.6 0.082 0.134
74 124.7 162.1 0.061 0.108 99 204.9 233.9 0.143 0.192
Table 2: Full evaluation of CIFAR100 on Federated CIFAR10.
E.5 Evaluate CIFAR100 on Federated ImageNet-32
We expand the evaluation of CIFAR100 to Federated ImageNet-32 dataset. Similarly, we extracted samples from each class of CIFAR-100 and treated them as the output of one hundred distinct generators, each corresponding to a single class. We also keep the first one hundred classes of ImageNet-32 and simulate one hundred clients. Each client hold all images ($\sim$1300) from a single class. We evaluate all the generators on Federated ImageNet-32 and the result is shown in Figure 16. The ranks provided by FID-avg and FID-all is inconsistent in a much more complex distributed learning setting.
Figure 16: Evaluate CIFAR100 on Federated ImageNet-32.
Appendix F Evaluation on Variance-Limited Federated Datasets
F.1 Experiment Setting
In the federated learning literature, it is relatively common that each client possesses only a small portion of the collective dataset, and the data diversity within each client’s holdings is significantly constrained. To illustrate, consider the case of smartphone users who exclusively own pictures of themselves, all of which share remarkable similarity. Nevertheless, in a network comprising millions of users, the overall dataset’s distribution still exhibits significant variance. In such scenarios, our theoretical framework suggests that the disparity between FID-all and FID-avg can become more pronounced. To experiment the effect of such distribution heterogeneity, we simulated and evaluated generative models under variance-limited federated datasets. To obtain a variance-limited federated dataset, for each class in the image dataset, we kept only a single image and its K-nearest neighbors. To find the $K$ nearest neighbors, we used the $L_{2}$-distance in the Inception-V3 2048-dimensional semantic space. It is worth noting that our experimentation has shown that varying K within the range of 5 to 100 does not alter the core conclusions. This approach effectively mimics scenarios where each client’s data has limited variance. We simulated the variance-limited federated learning setting for CIFAR-10, CIFAR-100 and a 32$\times$32 version of ImageNet (IN-32). For CIFAR-10 and CIFAR-100, we utilized all the classes in the dataset and for IN-32 we utilized the first 100 classes. We chose $K=20$ in the experiments. Intuitively, a larger $K$ leads to a more significant intra-client variance.
Variance-controlled Generators. To simulate a generator, we initiate the process by randomly selecting a sample from the dataset. We then gather its M-nearest neighbors from the original dataset (w/o federated learning setting). We consider this subset of samples as a set of generated samples generated by a generator denoted by $G_{M}$. By increasing the value of $M$, we generated a sequence of generators with progressively higher variance values. We tried the $M$ range from 100 to 50000. We evaluated all the generative models, denoted as $G_{M}$ with the chosen $M$ values, using the Variance-Limited Federated datasets.
F.2 Results on Variance-Limited Federated CIFAR10
The evaluation results on CIFAR10 are shown in Figure 17. Our findings reveal a distinct pattern in the behavior of FID-avg and FID-all as generator variance varies while the distance between KID-avg and KID-all remains the same. Our numerical results highlight the impact of the choice of FID-all and FID-avg on model rankings in federated learning settings with limited intra-client variance. , which can be broadly categorized into three phases.
F.3 Results on Variance-Limited Federated CIFAR100
Similar to the experiments on CIFAR10, we have also applied the variance-limited federated dataset setting to CIFAR100. We keeps K=20 images in each class. For variance-controlled generators, we select a sample from original CIFAR100 and gather the M-nearset neighbors. The range of M keeps the same with that in the previous subsection. We show the results in Figure 18. The results still support our main claims: FID-avg and FID-all gives inconsistent results while KID-avg and KID-all give the same.
Figure 17: Evaluation on Variance-Limited Federated CIFAR10.
Figure 18: Evaluation on Variance-Limited Federated CIFAR100.
F.4 Results on Variance-Limited Federated IN32
Results on ImageNet-32 are illustrated in Figure 19. We further plot the relationship between ranking given by FID-scores and KID-scores in Figure 20.
Figure 19: FID and KID-based Evaluations of variance-controlled generators on variance-limited federated ImageNet-32.
Figure 20: Comparing FID-based and KID-based rankings of variance-limited federated Mini-ImageNet-based simulated generative models. The lower the rank is, the better.
F.5 The Effect of Intra-Client Variance
In the main body of this paper, we choose $K=20$ when we conduct the variance-limited federated CIFAR10 dataset. Hyper-parameter K controls the intra-client variance, the larger the $K$ the larger the variance. The number of $K$ will not affect the key conclusion. We prove this claim by conducting an ablation study on hyper-parameter $K$. The $K$ is selected from {5,10,20,50} in our experiment. The results are illustrated in Figure 21. Each of these figure gives a U-shape curve, which indicates that the rankings given by FID-all and FID-avg are highly inconsistent, especially when the intra-client variance and inter-client variance are mismatched.
Figure 21: Ablation study on hyper-paramter K.
F.6 Precision/Recall and Density/Coverage Evaluations
Alongside FID and KID metrics, we expanded our analysis to include Precision/Recall [16] and Density/Coverage [17], adhering to the experimental frameworks outlined in Equation 2 and Equation 1.
Experiment Setup. For the Precision/Recall assessments, we employed the official implementation, setting the number of clusters to five, and evaluated the precision and recall metrics on a distributed CIFAR-10 dataset. In the case of Density/Coverage, we used the publicly available version of the evaluation metrics.
Results. As depicted in Figure 22, the precision and recall evaluations revealed variability in the consistency of rankings among generative models within a heterogeneous data environment. Specifically, while recall measurements (Recall-all and Recall-avg) consistently yielded identical values, the precision scores varied. This inconsistency in precision scores highlights potential challenges in using precision to rank models under distributed data distributions. Regarding Density and Coverage, our findings indicate that although density-based metrics (Density-all and Density-avg) sometimes provided inconsistent rankings, coverage-based assessments paralleled the recall metrics in consistently ranking models across evaluations.
Figure 22: Precision / Recall and Density / Coverage on distributed CIFAR10.